OpenStudy (anonymous):

Which graph represents the function f (x − 3), if f (x) = x2?

OpenStudy (anonymous):

could you draw it for me bro

OpenStudy (anonymous):

well it was wrong in any case

OpenStudy (anonymous):

\[f(x)=x^2\\ f(x-3)=(x-3)^2\]

OpenStudy (mathmale):

Given the graph of a function f(x), the graph of f(x-h) can be found simply by translating the entire graph of f(x) h units to the right. Try it. Graph y=x^2. Then graph y-(x-3)^2. Here, h=3.

OpenStudy (anonymous):

draw it for me cause the numers arent listed on the grraph

OpenStudy (anonymous):

here is your graph of \[(x-3)^{2}\]

OpenStudy (mathmale):

There are a couple of basic functions whose graphs you're really expected to know. that of y=x^2 is one of them. Please make up a reference page with the graphs of y=c y=x y=x^2 y=x^3 y=|x| and so on and learn them. In this particular case y=x^2 has a parabolic graph:

OpenStudy (mathmale):

The graph of y=(x-3)^2 looks exactly the same, EXCEPT that you must move the graph of y=x^2 3 units to the right.